# Three meta-questions on infinite-domain Constraint Satisfaction Problems

### v1

@@ A.1. Omitted definitions
For the other direction, we use a characterization via the pseudo-Siggers identity from [7, Theorems 1.3 and 3.4]. Suppose that $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$. Let $\mathfrak{C}$ be the model-complete core of $\mathfrak{A}$. Since $\mathfrak{A}$ and $\mathfrak{C}$ are homomorphically equivalent, $\mathfrak{C}$ also does not pp-construct $\mathfrak{K}_{3}$. Then, by Theorem 1.3 in [7], there exist $\alpha,\beta,s\in\mathrm{Pol}(\mathfrak{C})$ witnessing the pseudo-Siggers identity. Consider the structure $\mathbb{Q}^{\mathfrak{C}}$. By item 1, $\mathbb{Q}^{\mathfrak{C}}$ is an $\omega$-categorical model-complete core. Note that we can lift the homomorphic equivalence between $\mathfrak{A}$ and $\mathfrak{C}$ to a homomorphic equivalence between $\mathbb{Q}^{\mathfrak{A}}$ and $\mathbb{Q}^{\mathfrak{C}}$ by setting, for any homomorphism $h$ between $\mathfrak{A}$ and $\mathfrak{C}$ and irrespectively of its direction, $h^{\ast}(x,y)\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(e(x,y),h(y))$, where $e$ is an arbitrary injection from the Cartesian product of $\mathbb{Q}$ with the domain of $h$ into $\mathbb{Q}$. The inclusion of the injection in the first coordinate ensures that $h^{\ast}$ preserves $I_{4}$ and $\neq$ (because they are preserved by all injective operations). Since $\mathbb{Q}^{\mathfrak{A}}$ and $\mathbb{Q}^{\mathfrak{C}}$ are homomorphically equivalent, $\mathbb{Q}^{\mathfrak{C}}$ is the model-complete core of $\mathbb{Q}^{\mathfrak{A}}$. Next, we lift the pseudo-Siggers identity from $\mathrm{Pol}(\mathfrak{C})$ to $\mathrm{Pol}(\mathbb{Q}^{\mathfrak{C}})$. Let $e$ be an arbitrary injection from ×$\mathbb{Q}\times C$ into $\mathbb{Q}$. The structure $(\mathbb{Q};\neq,I_{4})$ is an $\omega$-categorical model-complete core that does not pp-construct $\mathfrak{K}_{3}$ [16, Theorems 12.0.1, 12.7.3, 12.9.2, and Corollary 6.4.4]. Hence, by Theorem 1.3 in [7] (since pp-constructions are more general than pp-interpretations with parameters), there are $\alpha^{\prime},\beta^{\prime},s^{\prime}\in\mathrm{Pol}(\mathbb{Q};\neq,I_{4})$ witnessing the pseudo-Siggers identity. We set
$\alpha^{\ast}(x,y)\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=$ $(\alpha^{\prime}(x),\alpha(y)),$
$\beta^{\ast}(x,y)\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=$ $(\beta^{\prime}(x),\beta(y)),$
$s^{\ast}((x_{1},y_{1}),\dots,(x_{6},y_{6}))\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=$ $(s^{\prime}(e(x_{1},y_{1}),\dots,e(x_{6},y_{6})),s(y_{1},\dots,y_{6})).$
@@
It is easy to verify that $\alpha^{\ast},\beta^{\ast},s^{\ast}\in\mathrm{Pol}(\mathbb{Q}^{\mathfrak{C}})$ and that these operations witness the pseudo-Siggers identity. It can be seen in the proof of Proposition 4.5(3) that also the structure $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ has less than double exponential orbit growth, and therefore so does its model-complete core $\mathbb{Q}^{\mathfrak{C}}$ by the proof of [7, Corollary 1.8]. Combining this with the pseudo-Siggers identity, it follows from Theorem 3.4 in [7] (whose item (ii) is equivalent to the pp-construction of $\mathfrak{K}_{3}$ by Theorem 1.8 in [9]) that $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$. ∎
A.7. A proof of Corollary 2.5
@@
Suppose that $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$. By Theorem 2.4(4), the structure $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$ either. Let $\mathfrak{C}$ be the model-complete core of $\mathfrak{A}$. By the proof of Theorem 2.4(4), $\mathbb{Q}^{\mathfrak{C}}_{I_{4}}$ is the model-complete core of $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$. Hence, also $\mathbb{Q}^{\mathfrak{C}}_{I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$. By [57, Theorem 3.14], $\mathrm{Pol}(\mathbb{Q}^{\mathfrak{C}}_{I_{4}})$ contains a binary essential operation $f$. Let $g$ and $h$ be any homomorphisms witnessing the homomorphic equivalence between $\mathbb{Q}^{\mathfrak{A}}_{I_{4}}$ and $\mathbb{Q}^{\mathfrak{C}}_{I_{4}}$. Then $(x,y)\mapsto g\circ f(h(x),h(y))$ is a binary essential polymorphism of $\mathbb{Q}^{\mathfrak{C}}_{I_{4}}$. Since $\mathbb{Q}^{\mathfrak{C}}_{I_{4}}$ is Pol-injective, this operation must be injective. ∎
Appendix B $\omega$-categorical sandwiches for PCSPs
@@
For $i\in[d]$, we denote by $\iota_{i}$ the function from $[1]$ to $[d]$ defined by $\iota_{i}(1)\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=i$.
@@
For (2), we proceed as in the proof of [16, Lemma 3.5.4]; the identity map yields a $d$-dimensional pp-interpretation of ${\mathfrak{A}}^{[d]}$ in $\mathfrak{A}$ and the first projection gives a 1-dimensional pp-interpretation of $\mathfrak{A}$ in ${\mathfrak{A}}^{[d]}.$
@@
For (4), let $\alpha:\mathfrak{D}_{1}\to\mathfrak{D}_{2}$ be an isomorphism between two finite substructures of ${\mathfrak{A}}^{[d]}$ and let $A_{1},A_{2}$ be the sets of elements of $A$ appearing as entries in some $d_{1}\in D_{1}$ or $d_{2}\in D_{2}$, respectively. Then the relations $S^{{\mathfrak{A}}^{[d]}}_{\iota_{i},\iota_{j}}$ and $\widehat{R}^{\,{\mathfrak{A}}^{[d]}}$ for $R\in\tau$ guarantee that $\alpha$ induces an isomorphism $\beta$ between $\mathfrak{A}_{1}$ and $\mathfrak{A}_{2}$. Since $\mathfrak{A}$ is homogeneous, we can extend $\beta$ to some automorphism $\beta^{\ast}$ of $\mathfrak{A}$. Now $\alpha^{\ast}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\beta^{\ast}\curvearrowright A^{d}$ is an automorphism of ${\mathfrak{A}}^{[d]}$ extending $\alpha$.
For (5), let $e\in\mathrm{End}({\mathfrak{A}}^{[d]})$ be arbitrary and let $D$ be an arbitrary finite subset of $A^{d}$. Define $A_{D}$ as the set of all elements of $A$ appearing in an entry of some $d\in D$. By item (1), we have $e=e^{\prime}\curvearrowright A^{d}$ for some $e^{\prime}\in\mathrm{End}(\mathfrak{A})$. Since $\mathfrak{A}$ is a model-complete core, there exists $\alpha^{\prime}\in\mathrm{Aut}(\mathfrak{A})$ such that $\alpha^{\prime}|_{A_{D}}=e^{\prime}|_{A_{D}}$. Then $\alpha\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\alpha^{\prime}\curvearrowright A^{d}$ is an automorphism of ${\mathfrak{A}}^{[d]}$ satisfying $\alpha|_{D}=e|_{D}$.
Item (6) can be shown in a similar fashion as [16, Lemma 3.5.4].
For (7), note that, by item (4), the mapping $\alpha\mapsto\alpha\curvearrowright A^{d}$ is a topological isomorphism from $\mathrm{Aut}(\mathfrak{A})$ to $\mathrm{Aut}({\mathfrak{A}}^{[d]})$. By Theorem A.1, we have that ${\mathfrak{A}}^{[d]}$ is homogeneous Ramsey.
@@
We split the proof of Proposition 5.2 into several individual statements, which are either independent or build on top of each other (in the order in which they are presented).
We start with the following lemma, a crucial independent observation about pseudo-cyclic polymorphisms from finite substructures of $\mathfrak{A}$ to $\mathfrak{A}$ itself.
@@
for all $s_{1},\dots,s_{n}\in S$. Suppose, on the contrary, that $f$ is not cyclic. Then there are $s_{1},\dots,s_{n}\in S$ such that $f(s_{1},\dots,s_{n})\neq f(s_{2},\dots,s_{n},s_{1})$. Let $<$ denote the linear order of $\mathfrak{B}$ and set $\tilde{s_{i}}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(s_{i},\dots,s_{n},s_{1},\dots,s_{i-1})$. By assumption, we have $f(\tilde{s_{1}})\neq f(\tilde{s_{2}})$. Without loss of generality, $f(\tilde{s_{1}})<f(\tilde{s_{2}})$. Since $\alpha_{1},\alpha_{2}$ preserve $<$ and its complement, we have
@@
The next lemma can be proved using a standard compactness argument.
Lemma B.2.
Let $\mathfrak{A}$ be an $\omega$-categorical structure and $R$ a finitary relation over $A$. If $\mathrm{Pol}(\mathfrak{A})$ preserves $R$, then, for every $n\in\mathbb{N}$, there exists a finite substructure $\mathfrak{S}$ of $\mathfrak{A}$ with $|S|\geq n$ such that $\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ preserves $R$.
Proof.
We need the following strengthening of [23, Lemma 10] tailored to our statement.
Claim B.3.
If $R$ is a union of $\ell$-many $\mathrm{Aut}(\mathfrak{A})$-orbits for some $\ell\in\mathbb{N}$ and there exists $g\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ that does not preserve $R$ for a finite substructure $\mathfrak{S}$ of $\mathfrak{A}$, then there exists an $\ell$-ary $g^{\prime}\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ that does not preserve $R$.
Proof.
By assumption, there exist $\bar{t}_{1},\dots,\bar{t}_{m}\in R$ such that $g(\bar{t}_{1},\dots,\bar{t}_{m})\notin R$. First, suppose that $m\leq\ell$. Define the $\ell$-ary operation $g^{\prime}$ by $g^{\prime}(x_{1},\dots,x_{\ell})\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=g(x_{1},\dots,x_{m})$. Clearly, $g^{\prime}\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$, and $g^{\prime}(\bar{t}_{1},\dots,\bar{t}_{m},\bar{t}_{m},\dots,\bar{t}_{m})=g(\bar{t}_{1},\dots,\bar{t}_{m})\notin R$. Next, suppose that $\ell\leq m$. Since $R$ is the union of $\ell$ orbits, we can pick $\ell$ representatives $\bar{t}_{i_{1}},\dots,\bar{t}_{i_{\ell}}$ of said orbits among $\bar{t}_{1},\dots,\bar{t}_{m}$. Without loss of generality, we may assume that $i_{j}=j$ for every $j\leq\ell$. Then, for every $\ell+1\leq i\leq m$, there exists $j_{i}\leq\ell$ and $\alpha_{i}\in\mathrm{Aut}(\mathfrak{A})$ such that $\bar{t}_{i}=\alpha_{i}(\bar{t}_{j_{i}})$. Define the $\ell$-ary operation $g^{\prime}$ by
$g^{\prime}(x_{1},\dots,x_{\ell})\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=g(x_{1},\dots,x_{\ell},\alpha_{\ell+1}(x_{j_{\ell+1}}),\dots,\alpha_{m}(x_{j_{m}})).$
Then $g^{\prime}\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ and $g^{\prime}$ does not preserve $R$ as witnessed by the tuple $\bar{t}_{1},\dots,\bar{t}_{\ell}$. ∎
Now we prove the statement of the lemma. Since $\mathfrak{A}$ is $\omega$-categorical, $R$ is a union of $\ell$ orbits for some $\ell\in\mathbb{N}$. Suppose, on the contrary, that there exists $n\in\mathbb{N}$ such that, for every finite substructure $\mathfrak{S}$ of $\mathfrak{A}$ with $|S|\geq n$, there exists $f_{S}\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ that does not preserve $R$. Then, by Claim B.3, for every $\mathfrak{S}$ with $|S|\geq n$ there also exists an $\ell$-ary $f^{\prime}_{S}\in\mathrm{Pol}(\mathfrak{S},\mathfrak{A})$ that does not preserve $R$. By a standard compactness argument (an application of Kőnig’s Tree Lemma), there is an $\ell$-ary polymorphism of $\mathfrak{A}$ that does not preserve $R$, a contradiction. ∎
The next lemma essentially states that $\mathcal{I}(\mathfrak{X})\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}={\mathfrak{X}}^{[d]}$ is a reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}({\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$. We will, however, use it in the more general setting of PCSPs.
@@
where $\mathrm{id}$ denotes the identity map on $[k]$. Thus, and since $h$ is a homomorphism, we have
@@
$\bar{y_{1}}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(y_{1},\dots,y_{i-1},x_{1},y_{i+1},\dots,y_{k}),\quad\bar{z_{1}}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(y_{1},\dots,y_{i-1},x_{1},y_{i+1},\dots,y_{k},x_{2},\dots,x_{2}),$
$\bar{y_{2}}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(y_{1},\dots,y_{i-1},x_{2},y_{i+1},\dots,y_{k}),\quad\bar{z_{2}}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(y_{1},\dots,y_{i-1},x_{2},y_{i+1},\dots,y_{k},x_{1},\dots,x_{1}).$
@@
Finally, we combine the two auxiliary constructions from before by setting $\tilde{\mathfrak{X}}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=\mathfrak{X}^{\prime}/{\sim}$.
@@
Suppose, on the contrary, that this is not the case. Then there is some bound $\mathfrak{F}$ of $\mathfrak{B}$ that embeds into $\tilde{\mathfrak{X}}$. Without loss of generality, $F=\{[x_{1}]_{\sim},\dots,[x_{n}]_{\sim}\}\subseteq\tilde{X}$. Since all bounds have size $\leq d$, we have $n\leq d$. As before, we shall use the abbreviations $\bar{x}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(x_{1},\dots,x_{d})$ and $\bar{a}\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=(a_{1},\dots,a_{d})$. Now, for every $k$-ary relation $R^{\tilde{\mathfrak{X}}}$ and every $\iota\colon[k]\to[n]$, we have
@@
By fixing some $x_{n+1},\dots,x_{d}\in X$ and setting $e([x_{i}]_{\sim})\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=a_{i}$ for $i\leq n$ and some $\bar{a}\in h(\bar{x})$, we have that $e$ is an embedding of $\mathfrak{F}$ into $\mathfrak{B}$, a contradiction. ∎
Let $e^{\prime}$ be an embedding from $\tilde{\mathfrak{X}}$ to $\mathfrak{B}$, which exists by Claim B.6. It follows from Claim B.5 that $e^{\prime}$ induces a homomorphism $g\colon\mathfrak{X}^{\prime}\to\mathfrak{B}$ via $g(x)\mathrel{\mathop{:}}\mathrel{\mkern-1.2mu}=e^{\prime}([x]_{\sim})$.
@@
Let $R\in\tau$ and $(x_{1},\dots,x_{k})\in R^{\mathfrak{X}}$ be arbitrary. Since $h$ is a homomorphism, we have $h(x_{1},\dots,x_{d})\in R_{\mathrm{id}}^{\,\smash{{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}}$ for all $x_{k+1},\dots,x_{d}\in X$. Since every automorphism of $\mathfrak{B}$ is also an automorphism of $\mathfrak{A}$, if follows from Claim B.7 that $(g(x_{1}),\dots,g(x_{k}))\in R^{\mathfrak{A}}$. ∎
@@
It is not hard to show that ${(\mathfrak{S}^{n})}^{[d]}\simeq({\mathfrak{S}}^{[d]})^{n}$ (cf. [59, Corollary 22]) for all $d,n\in\mathbb{N}$. Hence, every homomorphism $h\colon({\mathfrak{S}}^{[d]})^{n}\to{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$ is also a homomorphism from ${(\mathfrak{S}^{n})}^{[d]}$ to ${\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Suppose that $h$ is a function symbol appearing in $\Sigma$ that is assigned an element of $\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$. By Lemma B.4, there is a homomorphism $g\colon\mathfrak{S}^{n}\to\mathfrak{A}$ such that

### v2

@@ A.1. Omitted definitions
For the other direction, we use a characterization via the pseudo-Siggers identity from [7, Theorems 1.3 and 3.4]. Suppose that $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$. Let $\mathfrak{C}$ be the model-complete core of $\mathfrak{A}$. Since $\mathfrak{A}$ and $\mathfrak{C}$ are homomorphically equivalent, $\mathfrak{C}$ also does not pp-construct $\mathfrak{K}_{3}$. Then, by Theorem 1.3 in [7], there exist $\alpha,\beta,s\in\mathrm{Pol}(\mathfrak{C})$ witnessing the pseudo-Siggers identity. Consider the structure $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$. By item 1, $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ is an $\omega$-categorical model-complete core. Note that we can lift the homomorphic equivalence between $\mathfrak{A}$ and $\mathfrak{C}$ to a homomorphic equivalence between $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ by setting, for any homomorphism $h$ between $\mathfrak{A}$ and $\mathfrak{C}$ and irrespectively of its direction, $h^{\ast}(x,y)\coloneqq(e(x,y),h(y))$, where $e$ is an arbitrary injection from the Cartesian product of $\mathbb{Q}$ with the domain of $h$ into $\mathbb{Q}$. The inclusion of the injection in the first coordinate ensures that $h^{\ast}$ preserves $I_{4}$ and $\neq$ (because they are preserved by all injective operations). Since $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$ and $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ are homomorphically equivalent, $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ is the model-complete core of $(\mathbb{Q}\wr\mathfrak{A})_{\neq}$. Next, we lift the pseudo-Siggers identity from $\mathrm{Pol}(\mathfrak{C})$ to $\mathrm{Pol}((\mathbb{Q}\wr\mathfrak{C})_{\neq})$. Let $e$ be an arbitrary injection from ×$\mathbb{Q}\times C$ into $\mathbb{Q}$. The structure $(\mathbb{Q};\neq,I_{4})$ is an $\omega$-categorical model-complete core that does not pp-construct $\mathfrak{K}_{3}$ [15, Theorems 12.0.1, 12.7.3, 12.9.2, and Corollary 6.4.4]. Hence, by Theorem 1.3 in [7] (since pp-constructions are more general than pp-interpretations with parameters), there are $\alpha^{\prime},\beta^{\prime},s^{\prime}\in\mathrm{Pol}(\mathbb{Q};\neq,I_{4})$ witnessing the pseudo-Siggers identity. We set $\alpha^{\ast}(x,y)\coloneqq(\alpha^{\prime}(x),\alpha(y))$, $\beta^{\ast}(x,y)\coloneqq(\beta^{\prime}(x),\beta(y))$, and
$s^{\ast}((x_{1},y_{1}),\dots,(x_{6},y_{6}))\coloneqq(s^{\prime}(e(x_{1},y_{1}),\dots,e(x_{6},y_{6})),s(y_{1},\dots,y_{6})).$
@@
It is easy to verify that $\alpha^{\ast},\beta^{\ast},s^{\ast}$ are polymorphisms of $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ and that these operations witness the pseudo-Siggers identity. It moreover follows by inspection of the proof of Proposition 4.5(3) that also the structure $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ has less than double exponential orbit growth, and therefore so does its model-complete core $(\mathbb{Q}\wr\mathfrak{C})_{\neq}$ by the proof of [7, Corollary 1.8]. Combining this with the pseudo-Siggers identity, it follows from Theorem 3.4 in [7] (whose item (ii) is equivalent to the pp-construction of $\mathfrak{K}_{3}$ by Theorem 1.8 in [9]) that $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$. ∎
A.8. A proof of Corollary 2.5
@@
Suppose that $\mathfrak{A}$ does not pp-construct $\mathfrak{K}_{3}$. By Theorem 2.4(4), the structure $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$ either. Let $\mathfrak{C}$ be the model-complete core of $\mathfrak{A}$. By the proof of Theorem 2.4(4), $(\mathbb{Q}\wr\mathfrak{C})_{\neq,I_{4}}$ is the model-complete core of $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$. Hence, also $(\mathbb{Q}\wr\mathfrak{C})_{\neq,I_{4}}$ does not pp-construct $\mathfrak{K}_{3}$. By [55, Theorem 3.14], $\mathrm{Pol}((\mathbb{Q}\wr\mathfrak{C})_{\neq,I_{4}})$ contains a binary essential operation $f$. Let $g$ and $h$ be any homomorphisms witnessing the homomorphic equivalence between $(\mathbb{Q}\wr\mathfrak{A})_{\neq,I_{4}}$ and $(\mathbb{Q}\wr\mathfrak{C})_{\neq,I_{4}}$. Then $(x,y)\mapsto g\circ f(h(x),h(y))$ is a binary essential polymorphism of $(\mathbb{Q}\wr\mathfrak{C})_{\neq,I_{4}}$. Since $(\mathbb{Q}\wr\mathfrak{C})_{\neq,I_{4}}$ is Pol-injective, this operation must be injective. ∎
Appendix B $\omega$-categorical cheeses for PCSPs
@@
For $i\in[d]$, we denote by $\iota_{i}$ the function from $[1]$ to $[d]$ defined by $\iota_{i}(1)\coloneqq i$.
@@
For (2), we proceed as in the proof of [15, Lemma 3.5.4]; the identity map yields a $d$-dimensional pp-interpretation of ${\mathfrak{A}}^{[d]}$ in $\mathfrak{A}$ and the first projection gives a 1-dimensional pp-interpretation of $\mathfrak{A}$ in ${\mathfrak{A}}^{[d]}.$
@@
For (4), let $\alpha:\mathfrak{D}_{1}\to\mathfrak{D}_{2}$ be an isomorphism between two finite substructures of ${\mathfrak{A}}^{[d]}$ and let $A_{1},A_{2}$ be the sets of elements of $A$ appearing as entries in some $d_{1}\in D_{1}$ or $d_{2}\in D_{2}$, respectively. Then the relations $S^{{\mathfrak{A}}^{[d]}}_{\iota_{i},\iota_{j}}$ and $\widehat{R}^{\,{\mathfrak{A}}^{[d]}}$ for $R\in\tau$ guarantee that $\alpha$ induces an isomorphism $\beta$ between $\mathfrak{A}_{1}$ and $\mathfrak{A}_{2}$. Since $\mathfrak{A}$ is homogeneous, we can extend $\beta$ to some automorphism $\beta^{\ast}$ of $\mathfrak{A}$. Now $\alpha^{\ast}\coloneqq\beta^{\ast}\curvearrowright A^{d}$ is an automorphism of ${\mathfrak{A}}^{[d]}$ extending $\alpha$.
For (5), let $e\in\mathrm{End}({\mathfrak{A}}^{[d]})$ be arbitrary and let $D$ be an arbitrary finite subset of $A^{d}$. Define $A_{D}$ as the set of all elements of $A$ appearing in an entry of some $d\in D$. By item (1), we have $e=e^{\prime}\curvearrowright A^{d}$ for some $e^{\prime}\in\mathrm{End}(\mathfrak{A})$. Since $\mathfrak{A}$ is a model-complete core, there exists $\alpha^{\prime}\in\mathrm{Aut}(\mathfrak{A})$ such that $\alpha^{\prime}|_{A_{D}}=e^{\prime}|_{A_{D}}$. Then $\alpha\coloneqq\alpha^{\prime}\curvearrowright A^{d}$ is an automorphism of ${\mathfrak{A}}^{[d]}$ satisfying $\alpha|_{D}=e|_{D}$.
Item (6) can be shown in a similar fashion as [15, Lemma 3.5.4].
For (7), note that, by item (1) and since the topology of pointwise convergence is determined by the behaviour of functions on finite subsets, the mapping $\alpha\mapsto\alpha\curvearrowright A^{d}$ is a topological isomorphism from $\mathrm{Aut}(\mathfrak{A})$ to $\mathrm{Aut}({\mathfrak{A}}^{[d]})$. By Theorem A.1, we have that ${\mathfrak{A}}^{[d]}$ is homogeneous Ramsey.
@@
We split the proof of Proposition 5.2 into several individual statements, which are either independent or build on top of each other (in the order in which they are presented). We start with the following lemma, a crucial independent observation about pseudo-cyclic polymorphisms from finite substructures of $\mathfrak{A}$ to $\mathfrak{A}$ itself.
@@
for all $s_{1},\dots,s_{n}\in S$. Suppose, on the contrary, that $f$ is not cyclic. Then there are $s_{1},\dots,s_{n}\in S$ such that $f(s_{1},\dots,s_{n})\neq f(s_{2},\dots,s_{n},s_{1})$. Let $<$ denote the linear order of $\mathfrak{B}$ and set $\tilde{s_{i}}\coloneqq(s_{i},\dots,s_{n},s_{1},\dots,s_{i-1})$. By assumption, we have $f(\tilde{s_{1}})\neq f(\tilde{s_{2}})$. Without loss of generality, $f(\tilde{s_{1}})<f(\tilde{s_{2}})$. Since $\alpha_{1},\alpha_{2}$ preserve $<$ and its complement, we have
@@
The next lemma essentially states that $\mathcal{I}(\mathfrak{X})\coloneqq{\mathfrak{X}}^{[d]}$ is a reduction from $\mathrm{CSP}(\mathfrak{A})$ to $\mathrm{CSP}({\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$. We will, however, use it in the more general setting of PCSPs.
@@
where $\mathrm{id}$ denotes the identity map on $[k]$. This and the fact that $h$ is a homomorphism yield
@@
$\bar{y_{1}}\coloneqq(y_{1},\dots,y_{i-1},x_{1},y_{i+1},\dots,y_{k}),\quad\bar{z_{1}}\coloneqq(y_{1},\dots,y_{i-1},x_{1},y_{i+1},\dots,y_{k},x_{2},\dots,x_{2}),$
$\bar{y_{2}}\coloneqq(y_{1},\dots,y_{i-1},x_{2},y_{i+1},\dots,y_{k}),\quad\bar{z_{2}}\coloneqq(y_{1},\dots,y_{i-1},x_{2},y_{i+1},\dots,y_{k},x_{1},\dots,x_{1}).$
@@
Finally, we combine the two auxiliary constructions from before by setting $\tilde{\mathfrak{X}}\coloneqq\mathfrak{X}^{\prime}/{\sim}$.
@@
Suppose, on the contrary, that this is not the case. Then there is some bound $\mathfrak{F}$ of $\mathfrak{B}$ that embeds into $\tilde{\mathfrak{X}}$. Without loss of generality, $F=\{[x_{1}]_{\sim},\dots,[x_{n}]_{\sim}\}\subseteq\tilde{X}$. Since all bounds have size $\leq d$, we have $n\leq d$. As before, we shall use the abbreviations $\bar{x}\coloneqq(x_{1},\dots,x_{d})$ and $\bar{a}\coloneqq(a_{1},\dots,a_{d})$. Now, for every $k$-ary relation $R^{\tilde{\mathfrak{X}}}$ and every $\iota\colon[k]\to[n]$, we have
@@
By fixing some $x_{n+1},\dots,x_{d}\in X$ and setting $e([x_{i}]_{\sim})\coloneqq a_{i}$ for $i\leq n$ and some $\bar{a}\in h(\bar{x})$, we have that $e$ is an embedding of $\mathfrak{F}$ into $\mathfrak{B}$, a contradiction. ∎
Let $e^{\prime}$ be an embedding from $\tilde{\mathfrak{X}}$ to $\mathfrak{B}$, which exists by Claim B.4. It follows from Claim B.3 that $e^{\prime}$ induces a homomorphism $g\colon\mathfrak{X}^{\prime}\to\mathfrak{B}$ via $g(x)\coloneqq e^{\prime}([x]_{\sim})$.
@@
Let $R\in\tau$ and $(x_{1},\dots,x_{k})\in R^{\mathfrak{X}}$ be arbitrary. Since $h$ is a homomorphism, we have $h(x_{1},\dots,x_{d})\in R_{\mathrm{id}}^{\,\smash{{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}}}$ for all $x_{k+1},\dots,x_{d}\in X$. Since every automorphism of $\mathfrak{B}$ is also an automorphism of $\mathfrak{A}$, it follows from Claim B.5 that $(g(x_{1}),\dots,g(x_{k}))\in R^{\mathfrak{A}}$. ∎
@@
It is not hard to show that ${(\mathfrak{S}^{n})}^{[d]}\simeq({\mathfrak{S}}^{[d]})^{n}$ (cf. [57, Corollary 22]) for all $d,n\in\mathbb{N}$. Hence, every homomorphism $h\colon({\mathfrak{S}}^{[d]})^{n}\to{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$ is also a homomorphism from ${(\mathfrak{S}^{n})}^{[d]}$ to ${\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}}$. Suppose that $h$ is a function symbol appearing in $\Sigma$ that is assigned an element of $\mathrm{Pol}({\mathfrak{S}}^{[d]},{\mathfrak{A}}^{[d]}_{\!{}_{\vphantom{/}\smash{/\mathrm{Aut}(\mathfrak{B})}}})$. By Lemma B.2, there is a homomorphism $g\colon\mathfrak{S}^{n}\to\mathfrak{A}$ such that